# HARNESSING Nth ROOT GATES FOR ENERGY STORAGE

### v1

@@ I Introduction
Advances in technology have enabled the construction and manipulation of microscopic systems, opening the door to develop experimental thermal machines which lie far from the thermodynamic limit [1, 2, 3, 4, 5, 6, 7, 8]. Quantum effects and phenomena, such the entanglement of particles, become extremely important in the description and modelling of these systems. These quantum phenomena can give advantages over classical systems in thermodynamic processes [8, 9, 10, 11, 12]. Quantum batteries represent an exciting application of quantum principles, with the potential to revolutionize energy storage for diverse scenarios. This kind of device has been implemented in multiple systems [13, 14, 15, 16, 17] and ergotropy has been the figure of merit used to quantify its effectiveness. [18, 19, 20]. Quantum batteries may be based on collective quantum phenomena [21, 13]. Hence, considering protocols using many-body interactions for these new technologies is of significant importance. Protocols inspired by quantum computing have been proposed for thermal machines, and the construction and testing of quantum many-body systems supporting the use of quantum logic gates is experimentally realisable using, among others, trapped ions [22, 23], Rydberg atoms [24], and superconducting qubits [25]. While still in their early stages of development, theoretical and experimental studies have demonstrated the feasibility and potential benefits of quantum batteries [21, 26, 19, 27, 13, 28]. Understanding their underlying principles and exploring practical implementations is aim of this work. Replacing full Pauli gate operations by a fractional step-wise protocol has allowed for investigating the presence of quantum friction, i.e. the generation of quantum coherence in the instantaneous energy eigenbasis under a non-permuting protocol [29, 30]. The experimental realization of this protocol has proven that quantum-friction induces a violation of the work fluctuation dissipation relation, certifying an additional genuine quantum effect [31].
Here, we propose a fractional step-wise protocol with multiple qubits and using Nth-root controlled-not logic gates (NRCGs) for charging and discharging a quantum battery. We investigate the protocol for initial quantum coherences and find that this leads to improved performances as compared to a thermal state.
@@ II Theoretical outline
We investigate various quantum thermodynamic protocols for charging a quantum battery. We express the scheme in a gate-based approach with quantum circuits of few qubit systems, initialized and driven by a gate sequence and we characterize the outcome of the protocol using the amount of ergotropy that has been generated.
@@ II.1 System and quantum circuits
We consider a system composed either by two (A and B) or by three (A, B, and C) qubits that interact with each other through the use of NRCGs (Fig. 1). These gates may generate entanglement between the component qubits. Qubits A and C are initialised each in a thermal state at temperatures $T_{A}$ and $T_{C}$, respectively, where $T_{A}>T_{C}>0$. The single-qubit thermal state is a Gibbs state defined as [32]
@@
where $j=A,B,C$. As reference systems, we consider the ones with each qubit prepared in a thermal state: here, there are no initial quantum coherences. We compare these with the systems in which qubit B is prepared in a pure state: then initial quantum coherence will carry through the qubits when a circuit is applied. Qubit B is initialised in a pure state as $\rho_{Pure}=\left|\psi\rangle\langle\psi\right|$, where $|\psi\rangle=\cos\left(\theta/2\right)|0\rangle+e^{i\phi}\sin\left(\theta/2\right)|1\rangle$ with a value of $\theta$ which ranges from $0\ to\ \pi$ while $\phi$ ranges from 00 to $2\pi$. For this investigation we will mainly focus on $\phi=0$ and $\phi=\pi$, though the full range of $\phi$ values will also be considered. Here $|0\rangle$ is the ground state and $|1\rangle$ is the excited state. In the paper, energies are given in units of $\epsilon_{2B}$, which is then set to 1 in all calculations.
@@
with $H_{j}$ given by eq. (2) and $j=A,B$ or $j=A,B,C$ for two and three qubits, respectively. We note that eq. (3) represents the Hamiltonian for the total system at any time, including at the point of measurement, except when NRCGs are applied, inducing interactions between qubits.
@@ II.2 CNOT and Nth CNOT Root Logic Gates
where the first subscript represents the control qubit and the second subscript the target qubit. This type of gate may generate entanglement between two qubits which adds a level of quantumness to the system. The NRCG is a method of partially applying a CNOT gate, it is a unitary operation given by [33],
$N]{CNOT_{A,B}}\ =\ \begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&s&p\\ 0&0&p&s\end{pmatrix},\;\;\;\;\;\;\sqrt[N]{CNOT_{B,A}}\ =\ \begin{pmatrix}1&0&0&0\\ 0&s&0&p\\ 0&0&0&0\\ 0&p&0&s\end{pmatrix}.$ (5)
Here $s=\frac{1}{2}+\frac{1}{2}e^{\frac{i\pi}{N}}$ and $p=\frac{1}{2}-\frac{1}{2}e^{\frac{i\pi}{N}}$.
A cycle is defined as N iterations of the basic circuit, corresponding to $M=N$ in Fig. 1(a). In the limit of large $N$, $N$ consecutive applications of a NRCG with the same control and target qubits could be seen as a trotterization of the CNOT gate, aiming at explicitly implement the gate as an adiabatic dynamics. In this sense, our protocols gives explicit access to intermediate states, e.g. allowing for the opportunity to use states with different degree of entanglement from the end result of the full CNOT gate.
Figure 1: (a) Circuit diagram for the three-qubit protocols: the full protocol includes $M$ iterations, each composed by two blocks of gates (blue squares). If $M=N$, the protocol is said to run over a ”cycle”. (b) The protocols considered, i.e. case 1, case 2, and case 3, differ by the set of gates within each ’Controlled gate’ block, as indicated. A two-qubit protocol emerges from the circuit in panel (a) by removing qubit C and the second Controlled-gate block.
@@ II.3 Circuit
A schematics of the protocols applied is shown in Fig. 1. We consider systems of two and three qubits and for each of them we examine three ways of applying interactions, labeled 1, 2 and 3 in panel (b). Case 1 and 2 examine the dynamics when one of the qubit is never the control qubit, being this either the qubit A or C for the three-qubit systems. Case 3 examines the protocol where all qubits are either a control or target qubit over the course of one iteration. In all protocols, qubits A and C are always initialised in a thermal state using eq. 1 whereas qubit B is initialised in either a pure or a thermal state. This allows the probing how initial quantum coherences affect the system’s capabilities as a battery.
The system is evolved, corresponding to the quantum circuits in Fig. 1 and to the evolution
@@
where $\rho^{(0)}_{sys}$ is the initial density matrix of the total system and $\rho^{(f)}_{sys}$ indicates the total density matrix after one iteration. The unitary $U$ represents the controlled gates specified in Fig. 1, as appropriate to each protocol, with NRCG’s of the form Eq. 5.
We note that in all three-qubit protocols, as NRCGs are applied to different qubits or alternating different control/target qubits, successive iterations will not generate full CNOT gates. Instead, for Cases 1 and 2 and two-qubit system, multiples of $N$ iterations of the circuit will be equivalent to the application of multiple standard CNOT gates; however, fractions of $N$ iterations will allow access to intermediate states in the evolution leading to a CNOT.
@@ II.4 Ergotropy
where $\{s_{j}\}$ are the eigenvalues of $\rho$ and $\{|j\rangle\}$ the eigenstates of $H$ [34]. No work can be extracted from a passive state, and for all unitaries $U$, $Tr\left(\sigma_{{\rho}}H\right)\leq Tr\left(U\sigma_{{\rho}}U^{\dagger}H\right)$ [34, 35].
@@ III Plan of the paper and anticipation of main results
The application of Nth-root gate operation between qubits unlocks a couple of assets, as we are able to follow in a step-wise protocol the time evolution and investigate the functioning of quantum battery protocols in mode detail. Consequently, we consider three different protocols highlighted in Fig. 1 for varying number of iterations and different initialization. For a fair and meaningful comparison among all these systems and protocols, we will examine four measures of performance for an optimal battery:
@@
  • Ergotropy variation, the difference between the final and initial ergotropies, relevant when the initial state, like here, is not a passive state;
  • The ratio between ergotropy and final energy of the system, which describes the fraction of energy that can be extracted as work
  • A figure of merit, combining the ergotropy variation and the ratio.
We find that results are strongly affected by the initial coherence of qubit B, and the same circuit may lead to minimal or to maximal ergotropy, depending on the setting of the phase. Also, the overall largest ergotropy variation and figure of merit are obtained from a circuit with high initial coherence. For all protocols and all measures of performance, initializing qubit B in a pure state as opposed to a thermal state provides substantial advantage.
@@ IV.1 Ergotropy scans for two- and three-qubit systems
The two-qubit system is initialised with qubit A thermalised in the Gibbs state eq. (1). We are using a reservoir where $k_{B}T_{A}=4\;\epsilon_{2B}$ For the three-qubit system qubits A and C are initialised at $k_{B}T_{A}=4\;\epsilon_{2B}$ and $\ k_{B}T_{C}=0.4\;\epsilon_{2B}$, respectively. Qubit B is initialised in the pure state $|\psi\rangle=\cos(\theta/2)|0\rangle+\exp(i\phi)\sin(\theta/2)|1\rangle$. One cycle is composed by $N$ iterations. We choose $N=15$, with this value showing a smooth evolution of the state of the system as iterations are applied. Cycles with $N>15$ shows no improvement but are increasing the run time of the simulations. Cycles with $N<15$ show increasing discontinuities, moving the evolution closer to the one using a complete CNOT gate ($N=1$). For details, see the appendix Fig 10.
Figure 2: Ergotropy of the total system for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ Iterations $\ \leq\ 30$ (y-axis) for two-qubit systems (first row) and three-qubit system (second row), where columns from left to right are case 1, 2, and 3 respectively. Brighter shades correspond to a greater value of ergotropy. Parameters are $\epsilon_{1j}=0\;\epsilon_{2B},\ \epsilon_{2j}=1\;\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$. $\phi=\pi$
@@
Each panel of Fig. 2 shows results for the ergotropy calculated after 1, 2, $\dots,~{}2N$ iterations, for $0\leq\theta\leq\pi$ and $\phi=\pi$. Increasing $\theta$ increases qubit B initial internal energy from $\epsilon_{1B}=0$ to $\epsilon_{2B}$. The first row refers to two-qubit systems, the second to results for three-qubit systems, for all three cases described in Fig. 1(b), as labelled. We refer to each of the panel as an ”ergotropy scan”.
The initial ergotropy (corresponding to zero iterations, top line of all panels) mainly depends on the initial energy of qubit B. The closer $\theta$ is to $\pi$ the larger the initial ergotropy of the total system. This can be seen across all cases for both two and three-qubit systems. However, how the battery behaves for increasing number of iterations depends on both the number of qubits and the circuit chosen, showing charging (ergotropy increasing with iterations) and discharging (ergotropy decreasing with iterations) regimes. For example, for case 1, $\theta=\pi$ and iterations increasing from 0 to 15, the two qubit system would discharge while the three qubit system would be charging.
Cases 1 and 3 show the largest difference in the ergotropy evolution, both with respect to the initial ergotropy value and when comparing the results for the two and three-qubit systems. It is clear that in cases 1 and 3 there is an optimal number of iterations that would maximise the ergotropy, with the region of largest ergotropy reached for three qubits, about 10 to 25 iterations and $\theta\gtrsim 1.9$ radians. For Case 2 we find only a small variation in ergotropy as the system evolves.
Results with $\phi=0$ are similar (see Appendix C, Fig. 13), with the three-qubit system for both case 1 and 3 showing most notable variation in ergotropy, though their maximum ergotropy region gets shifted towards larger values of $\theta$, $\theta\gtrsim 2.5$ radians, and they show a lesser variation of $W_{max}$ with iterations.
IV.2 Comparison between circuits with Full CNOT and NRCG
@@
Figure 3: Comparing the ergotropy $W_{max}$ for the NCRG-protocol with $0\ \leq$ N $\ \leq\ 30$ (magenta), and the standard CNOT (blue), in case of the two-qubit systems (first row) and three-qubit system (second row), where columns from left to right are cases 1, 2, and 3 respectively. Y-axis scale differs for each panel. The chosen value of $\theta$ is indicated in each panel. Parameters: $\Phi=\pi$, $\epsilon_{1j}=0\;\epsilon_{2B},\ \epsilon_{2j}=1\;\epsilon_{2B}$ for $j=A,B$, and $C$; $k_{B}T_{A}=4\;\epsilon_{2B}$ and $k_{B}T_{C}=0.4\;\epsilon_{2B}$. Magenta line: N = 15, Blue line: N = 1.

### v2

@@ I Introduction
Advances in technology have enabled the construction and manipulation of microscopic systems, opening the door to develop experimental thermal machines which lie far from the thermodynamic limit [1, 2, 3, 4, 5, 6, 7, 8]. Quantum effects and phenomena, such as the entanglement of particles, become extremely important in the description and modelling of these systems. These quantum phenomena can give advantages over classical systems in thermodynamic processes [8, 9, 10, 11, 12]. Quantum batteries represent an exciting application of quantum principles, with the potential to revolutionize energy storage for diverse scenarios. This kind of device has been implemented in multiple systems [13, 14, 15, 16, 17] and ergotropy has been the figure of merit used to quantify its effectiveness [18, 19, 20]. Quantum batteries may be based on collective quantum phenomena [21, 13]. Hence, considering protocols using many-body interactions for these new technologies is of significant importance. Protocols inspired by quantum computing have been proposed for thermal machines, and the construction and testing of quantum many-body systems supporting the use of quantum logic gates is experimentally realisable using, among others, trapped ions [4, 22], Rydberg atoms [23], and superconducting qubits [24]. While still in their early stages of development, theoretical and experimental studies have demonstrated the feasibility and potential benefits of quantum batteries [21, 25, 19, 26, 13, 14, 27]. Understanding their underlying principles and exploring practical implementations is the aim of this work. Replacing full Pauli gate operations with a fractional step-wise protocol has allowed for investigating the presence of quantum friction, i.e., the generation of quantum coherence in the instantaneous energy eigenbasis under a non-permuting protocol [28, 29]. The experimental realization of this protocol has proven that quantum friction induces a violation of the work fluctuation dissipation relation, certifying an additional genuine quantum effect [30].
Here, we propose a fractional step-wise protocol with multiple qubits and using experimentally realisable Nth-root controlled-not logic gates (NRCGs) for charging and discharging a quantum battery. Repeat applications of NRCGs act as a trotterization of controlled-not logic gates; this has some analogy with collision models used in other works involving the charging of quantum batteries [31, 32, 33, 22]. The NRCG stepped approach to the dynamics allows access to intermediate ergotropy states typically unobtainable using full controlled-not logic gates. This offers a greater control of the quantity of energy stored in the battery. NRCGs are feasible with current technology, and we find that, depending on the protocol and initial conditions, the use of an NRCG may provide overall more ergotropy than the use of its full controlled-not gate counterpart. We investigate the protocol for initial quantum coherences and find that this leads to improved performances as compared to a thermal state.
@@ II Theoretical outline
We investigate various quantum thermodynamic protocols for charging a quantum battery. We express the scheme in a gate-based approach with quantum circuits of few-qubit systems, initialised and driven by a gate sequence, where we characterize the outcome of the protocol using the amount of ergotropy that has been generated.
@@ II.1 System and Quantum Circuits
We consider a system composed either by two (A and B) or by three (A, B, and C) qubits that interact with each other through the use of NRCGs (Figure 1). These gates may generate entanglement between the component qubits. Qubits A and C are initialised each in a thermal state at temperatures $T_{A}$ and $T_{C}$, respectively, where $T_{A}>T_{C}>0$. The single-qubit thermal state is a Gibbs state defined as [34]
@@
where $j=A,B,C$. As reference systems, we consider the ones with each qubit prepared in a thermal state: here, there are no initial quantum coherences. We compare these with the systems in which qubit B is prepared in a pure state; then, initial quantum coherence will carry through the qubits when a circuit is applied. Qubit B is initialised in a pure state as $\rho_{Pure}=\left|\psi\rangle\langle\psi\right|$, where $|\psi\rangle=\cos\left(\theta/2\right)|0\rangle+e^{i\phi}\sin\left(\theta/2\right)|1\rangle$ with a value of $\theta$ which ranges from $0\ to\ \pi$, while $\phi$ ranges from 00 to $2\pi$. For this investigation, we will mainly focus on $\phi=0$ and $\phi=\pi$, though the full range of $\phi$ values will also be considered. Here, $|0\rangle$ is the ground state and $|1\rangle$ is the excited state. In this paper, energies are given in units of $\epsilon_{2B}$, which is then set to 1 in all calculations.
@@
with $H_{j}$ given by Equation (2) and $j=A,B$ or $j=A,B,C$ for two and three qubits, respectively. We note that Equation (3) represents the Hamiltonian for the total system at any time, including at the point of measurement, except when NRCGs are applied, inducing interactions between qubits.
@@ II.2 CNOT and Nth CNOT Root Logic Gates
where the first subscript represents the control qubit and the second subscript, the target qubit. This type of gate may generate entanglement between two qubits, which adds a level of quantumness to the system. Entanglement of intermediate states, which a CNOT gate may achieve, is known to be necessary to gain a quantum advantage in battery charging operations [35]. The NRCG is a method of partially applying a CNOT gate; it is a unitary operation given by [36],
$N]{CNOT_{A,B}}\ =\ \begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&s&p\\ 0&0&p&s\end{pmatrix},\;\;\;\;\;\;\sqrt[N]{CNOT_{B,A}}\ =\ \begin{pmatrix}1&0&0&0\\ 0&s&0&p\\ 0&0&1&0\\ 0&p&0&s\end{pmatrix}.$ (5)
Here, $s=\frac{1}{2}+\frac{1}{2}e^{\frac{i\pi}{N}}$ and $p=\frac{1}{2}-\frac{1}{2}e^{\frac{i\pi}{N}}$.
A cycle is defined as N iterations of the basic circuit, corresponding to $M=N$ in Figure 1a. In the limit of a large $N$, $N$ consecutive applications of an NRCG with the same control and target qubits could be seen as a trotterization of the CNOT gate, aiming at explicitly implementing the gate as an adiabatic dynamic. In this sense, our protocols give explicit access to intermediate states, e.g., allowing for the opportunity to use states with different degrees of entanglement from the end result of the full CNOT gate.
Figure 1: (a) Circuit diagram for the three-qubit protocols: the full protocol includes $M$ iterations, each composed of two blocks of gates (blue squares). If $M=N$, the protocol is said to run over a “cycle”. (b) The protocols considered, i.e., case 1, case 2, and case 3, differ by the set of gates within each ‘controlled gate’ block, as indicated. A two-qubit protocol emerges from the circuit in panel (a) by removing qubit C and the second controlled gate block.
@@ II.3 Circuit
A schematic of the protocols applied is shown in Figure 1. We consider systems of two and three qubits and for each of them, we examine three ways of applying interactions, labeled 1, 2, and 3, see Figure 1(b). Case 1 and 2 examine the dynamics when one of the qubits is never the control qubit, with this being either qubit A or C for the three-qubit systems. Case 3 examines the protocol where all qubits are either a control or target qubit over the course of one iteration. In all protocols, qubits A and C are always initialised in a thermal state using Equation (1), whereas qubit B is initialised in either a pure or a thermal state. This allows for probing how initial quantum coherences affect the system’s capabilities as a battery.
The system is evolved, corresponding to the quantum circuits in Figure 1 and to the evolution
@@
where $\rho^{(0)}_{sys}$ is the initial density matrix of the total system and $\rho^{(f)}_{sys}$ indicates the total density matrix after one iteration. The unitary $U$ represents the controlled gates specified in Figure 1, as appropriate to each protocol, with NRCGs of the form in Equation (5).
We note that in all three-qubit protocols, as NRCGs are applied to different qubits or alternate different control/target qubits, successive iterations will not generate full CNOT gates. Instead, for Cases 1 and 2 and the two-qubit system, multiples of $N$ iterations of the circuit will be equivalent to the application of multiple standard CNOT gates; however, fractions of $N$ iterations will allow access to intermediate states in the evolution leading to a CNOT.
@@ II.4 Ergotropy
where $\{s_{j}\}$ are the eigenvalues of $\rho$ and $\{|j\rangle\}$ are the eigenstates of $H$ [37]. No work can be extracted from a passive state, and for all unitaries $U$, $Tr\left(\sigma_{{\rho}}H\right)\leq Tr\left(U\sigma_{{\rho}}U^{\dagger}H\right)$ [37, 38].
@@ III Plan of the Paper and Anticipation of Main Results
The application of Nth-root gate operation between qubits unlocks a couple of assets, as we are able to follow the time evolution in a step-wise protocol and investigate the functioning of quantum battery protocols in more detail. Consequently, we consider three different protocols, highlighted in Figure 1 for varying numbers of iterations and different initializations. For a fair and meaningful comparison among all these systems and protocols, we will examine four measures of performance for an optimal battery:
@@
  • Ergotropy variation, which is the difference between the final and initial ergotropies, relevant when the initial state, like that here, is not a passive state;
  • The ratio between ergotropy and the final energy of the system, which describes the fraction of energy that can be extracted as work;
  • A figure of merit, which combines the ergotropy variation and the ratio.
We find that results are strongly affected by the initial coherence of qubit B, and the same circuit may lead to minimal or to maximal ergotropy, depending on the setting of the phase. Also, the overall largest ergotropy variation and figure of merit are obtained from a circuit with high initial coherence. For all protocols and all measures of performance, initializing qubit B in a pure state as opposed to a thermal state provides a substantial advantage.
@@ IV.1 Ergotropy Scans for Two- and Three-Qubit Systems
The two-qubit system is initialised with qubit A thermalised in the Gibbs state Equation (1). We are using a reservoir where $k_{B}T_{A}=4\epsilon_{2B}$ For the three-qubit system, qubits A and C are initialised at $k_{B}T_{A}=4\epsilon_{2B}$ and $\ k_{B}T_{C}=0.4\epsilon_{2B}$, respectively. Qubit B is initialised in the pure state $|\psi\rangle=\cos(\theta/2)|0\rangle+\exp(i\phi)\sin(\theta/2)|1\rangle$.
@@
One cycle is composed by $N$ iterations. We choose $N=15$, with this value showing a smooth evolution of the state of the system as iterations are applied. Cycles with $N>15$ show no improvement but are increasing the run times of the simulations. Cycles with $N<15$ show increasing discontinuities, moving the evolution closer to the one using a complete CNOT gate ($N=1$). For details, see Appendix A, Figure 10.
@@
Each panel of Figure 2 shows results for the ergotropy calculated after 1, 2, $\dots,~{}2N$ iterations, for $0\leq\theta\leq\pi$ and $\phi=\pi$. Increasing $\theta$ increases qubit B’s initial internal energy from $\epsilon_{1B}=0$ to $\epsilon_{2B}$. The first row refers to two-qubit systems, the second, to results for three-qubit systems, for all three cases described in Figure 1b, as labelled. We refer to each of the panels as an “ergotropy scan”.
@@
Figure 2: Ergotropy of the total system for $0\ \leq\ \theta\ \leq\ \pi$ (x-axis) and $0\ \leq$ Iterations $\ \leq\ 30$ (y-axis) for two-qubit systems (first row) and three-qubit systems (second row), where columns from left to right are case 1, 2, and 3, respectively. Brighter shades correspond to a greater value of ergotropy. Parameters are $\epsilon_{1j}=0\epsilon_{2B},\ \epsilon_{2j}=1\epsilon_{2B}$ for $j=A,B$, and $C$. $k_{B}T_{A}=4\epsilon_{2B}$ and $k_{B}T_{C}=0.4\epsilon_{2B}$. $\phi=\pi$.
The initial ergotropy (corresponding to zero iterations, top line of all panels) mainly depends on the initial energy of qubit B. The closer $\theta$ is to $\pi$, the larger the initial ergotropy of the total system. This can be seen across all cases for both two- and three-qubit systems. However, how the battery behaves for an increasing number of iterations depends on both the number of qubits and the circuit chosen, showing charging (ergotropy increasing with iterations) and discharging (ergotropy decreasing with iterations) regimes. For example, for case 1, $\theta=\pi$ and iterations increasing from 0 to 15, the two-qubit system would discharge while the three-qubit system would be charging.
Cases 1 and 3 show the largest difference in the ergotropy evolution, both with respect to the initial ergotropy value and when comparing the results for the two- and three-qubit systems. It is clear that in cases 1 and 3, there is an optimal number of iterations that would maximise the ergotropy, with the region of the largest ergotropy reached for three qubits, about 10 to 25 iterations and $\theta\gtrsim 1.9$ radians. For case 2, we find only a small variation in ergotropy as the system evolves.
Results with $\phi=0$ are similar (see Appendix C, Figure 13), with the three-qubit system for both case 1 and 3 showing the most notable variation in ergotropy, though their maximum ergotropy region gets shifted towards larger values of $\theta$, $\theta\gtrsim 2.5$ radians, and they show a lesser variation in $W_{max}$ with iterations.
IV.2 Comparison Between Circuits with Full CNOT and NRCGs